# Invert4TVG: A Temporal Video Grounding Framework with Inversion Tasks for Enhanced Action Understanding

### v1

@@ Abstract
Temporal Video Grounding (TVG) seeks to localize video segments matching a given textual query. Current methods, while optimizing for high temporal Intersection-over-Union (IoU), often overfit to this metric, compromising semantic action understanding in the video and query, a critical factor for robust TVG. To address this, we introduce Inversion Tasks for TVG (Invert4TVG), a novel framework that enhances both localization accuracy and action understanding without additional data. Our approach leverages three inversion tasks derived from existing TVG annotations: (1) Verb Completion, predicting masked action verbs in queries from video segments; (2) Action Recognition, identifying query-described actions; and (3) Video Description, generating descriptions of video segments that explicitly embed query-relevant actions.These tasks, integrated with TVG via a reinforcement learning framework with well-designed reward functions, ensure balanced optimization of localization and semantics. Experiments show our method outperforms state-of-the-art approaches, achieving a 7.1% improvement in R1@0.7 on Charades-STA for a 3B model compared to Time-R1. By inverting TVG to derive query-related actions from segments, our approach strengthens semantic understanding, significantly raising the ceiling of localization accuracy.
@@
Figure 1: Left side: A specific example of temporal video grounding. Through the model’s reasoning process, it can be observed that our method achieves a better understanding of actions in the video compared to Time-R1 (Xu 2025a) and other approaches. Right side: statistical results demonstrating that Time-R1, which is optimized solely for IoU loss, improves localization accuracy but reduces action understanding accuracy (where VC, AR, and VD are three Invert-TVG tasks proposed in this paper). By introducing Invert-TVG tasks, our method not only enhances action understanding but also boosts the accuracy of TVG (as shown in R1@0.3, R1@0.5, and R1@0.7).
Introduction
@@
Temporal Video Grounding (TVG) is crucial for long-form video understanding (Gaidon, Harchaoui, and Schmid 2013; Laptev and Pérez 2007; Darrell and Pentland 1993). It localizes a video segment matching a textual query (Gao et al. 2017; Zhang et al. 2023), enabling applications like video-text retrieval (Zhang 2024) and UAV positioning (Ju 2024).
@@
Existing TVG approaches fall into three paradigms: (1) Traditional methods using hand-crafted features, sliding windows, and DETR-like networks (Shi et al. 2022; Gordeev 2024); (2) LVLMs (Bai et al. 2025; Li et al. 2023) that regress segments via pretraining; and (3) RL-finetuned LVLMs, such as TimeZero (Xu 2025b) and Time-R1 (Xu 2025a), which use RL with format rewards for structured reasoning and IoU rewards for alignment.
TVG is fundamentally a video understanding task, with a particular emphasis on comprehension of actions, since textual queries typically contain verbs describing human activities (e.g., “opening a door”). These action-related verbs are critical for aligning video content with the query intent.
However, existing TVG methods primarily optimize for localization metrics like Intersection-over-Union (IoU), which measure the temporal overlap between predicted and ground-truth segments. However, models may overfit to superficial localization cues (e.g., relying on visual motion patterns rather than action semantics) and neglect deeper alignment between video actions and query verbs, leading to superficial matching rather than true understanding.
Figure 1 demonstrates this. Time-R1 (Xu 2025a), optimized for IoU in TVG, shows superior performance on TVG metrics (e.g., R1@0.3/R1@0.5/R1@0.7) compared to the baseline Qwen2.5-VL-3B. However, it exhibits degradation in key action-centric understanding tasks (VC/VD/AR), which ultimately hampering the TVG accuracy.
A key insight of our work is that improving action-level semantics is essential for pushing the upper bound of TVG accuracy. Conventional solutions involve joint training TVG tasks with other video understanding tasks such as VQA or video captioning. This is exactly how existing large vision-language models (LVLMs), e.g., Qwen2.5-VL series, are trained. However, this paradigm requires data expansion. More critically, VQA and captioning tasks are not originally designed to address action understanding in TVG scenarios. The research question in this paper is: Can we enhance TVG while preserving video-query action comprehension within existing TVG resources?
@@
We propose Invert4TVG, a novel framework that derives auxiliary action understanding tasks from TVG datasets without requiring external data or annotations. Specifically, we invert the input and output of the TVG task to obtain Invert-TVG tasks. Originally, given a video and a natural language query, TVG models predict the temporal segment where the action occurs. Inversely, given a video segment, Invert-TVG tasks infer the action-related information defined in the query. We introduce three Invert-TVG tasks: (1) Verb Completion (VC): Mask and infer action verbs from video segments. (2) Action Recognition (AR): Identify query actions from the segment. (3) Video Description (VD): Generate segment descriptions that contain action verbs in the queries.
With the well-defined Invert-TVG tasks, we then propose a reinforcement learning framework that executes TVG with high probability and samples Invert-TVG tasks equally, balancing localization and action understandings. Figure 1 demonstrates that our method, Invert4TVG, effectively preserves action understanding capabilities, thereby achieving superior performance on the TVG task compared to Time-R1, which solely optimizes the TVG objective without explicit action understanding. Our contributions include:
@@
  • We identify action understanding degradation in TVG from metric over-optimization, and address the problem via inversion tasks without extra data.
@@
  • We design Invert-TVG tasks which are self-supervised tasks repurposing TVG annotations for action understanding (including verb completion, action recognition, and video description).
@@
  • We propose a multi-task reinforcement learning framework dynamically balancing TVG and Invert-TVG, ensuring robust localization and understanding.
Related Works
@@
Temporal Video Grounding. Temporal Video Grounding (TVG) (Gao et al. 2017; Hendricks et al. 2017) localizes specific segments in untrimmed videos based on natural language queries. Recent methods fall into two categories: feature-based and frame-based LVLM approaches. Feature-based methods (Carreira and Zisserman 2017; Lin et al. 2022) extract video and text features using pretrained encoders, then predict timestamps via multimodal fusion. These rely heavily on feature quality, limiting performance. Frame-based LVLM methods (Li et al. 2024; Ren et al. 2024; Zeng et al. 2025) use supervised fine-tuning with autoregressive losses to generate event sequences, but often underperform on benchmarks like Charades-STA compared to feature-based approaches. To address this, Time-R1 (Xu 2025a) employs reinforcement learning (RL) with IoU rewards, achieving state-of-the-art TVG performance. However, its focus on temporal metrics neglects semantic alignment, constraining long-form video understanding
@@
RL in LVLMs. RL has advanced post-training of LVLMs through Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al. 2022; Yu et al. 2024) and Reinforcement Learning with Verifiable Reward (RLVR) (DeepSeek-AI 2025; Chen et al. 2025). RLHF aligns models with human preferences, improving tasks like image captioning, while RLVR enhances deterministic tasks like visual grounding (Liu et al. 2025). However, RL applications in long-form video tasks remain underexplored due to temporal complexity and semantic challenges. TimeZero (Xu 2025b), Time-R1 (Xu 2025a), VideoChat-R1 (Li et al. 2025) apply RL to TVG but overlooks semantic understanding degradation from IoU-focused rewards. Our Invert4TVG framework addresses this by repurposing TVG data into self-supervised tasks, enhancing action semantic understanding and surpassing traditional RL limitations in video grounding.
Method
@@
Figure 2: We propose three Invert-TVG tasks. By partially reversing the inputs and outputs of the TVG task we obtain Verb Completion, Action Recognition and Video Description, which reuse the original TVG dataset by taking ground truth video segments as input to reconstruct the target query related events. The prompts for the three invert-TVG tasks are not identical. For VC, the verb in the query is removed, and the model is required to complete and fill in this verb. AR asks the model to directly estimate the verb in the video. VD requires the model to describe the video content containing action verbs in the query.
@@
The TVG task aims to temporally localize video segments within long-form videos based on natural language queries. Given a video $V$, and a language query $q$, the goal is to identify the temporal boundaries $\tau=[t_{s},t_{e}]$ of the segment that best corresponds to $q$, where $t_{s},t_{e}\in\mathbb{R}^{+}$. The formal definition of the TVG task is as follows:
@@
Our approach is fundamentally a reinforcement learning algorithm that fine-tunes LVLMs (specifically the Qwen2.5-VL model series) by integrating both TVG and Invert-TVG tasks. In the following, we first introduce the fundamentals of GRPO (i.e., Group Relative Policy Optimization, a reinforcement learning algorithm proposed in (DeepSeek-AI 2025)). Next, we introduce the proposed inversion TVG tasks, together with reward functions used to train the TVG and Invert-TVG tasks. Finally, we introduce our Invert4TVG reinforcement learning framework.
Preliminary of Group Relative Policy Optimization
DeepSeek-R1 (DeepSeek-AI 2025), an early R1-style open-source LLM, uses GRPO to train policy $\pi_{\theta}$ for reasoning before answers, suitable for precise tasks like math. For query $q$, it generates responses $o_{1},\dots,o_{G}$} with score with $r(\cdot)$, and maximizes:
@@
Invert-TVG Tasks and Reward Functions
@@
In Temporal Video Grounding (TVG), the accuracy of model inference highly depends on its understanding of both the video $V$ and the query $q$. Therefore, we do not rely solely on the IoU reward but also introduce additional rewards to enhance or at least keep the model’s action understanding capability. To this end, as illustrated in Figure 2, we design three Invert-TVG tasks and define reward functions measuring how these tasks are fulfilled.
@@
Verb Completion. Verb Completion is a task that masks verbs in the query and asks the model to recover the verbs from the ground truth video segment. For example, if the original query is “Person closed the door”, the masked sentence is “Person [ ] the door”. The prompt is “Add a verb for the event ‘Person [ ] the door’ based on the video”. As long as the model outputs a sentence successfully recovering the verbs of the ground truth, a reward is given. For instance, if the output sentence is “Person [closes] the door” a full reward is given. Due to the randomness in the model’s output, we are primarily concerned with whether the model comprehends the actions within the relevant segments. Therefore, we design the system to treat verbs in different tenses as equivalent. The reward function is as follows:
$\mathrm{r}_{\mathrm{VC}}(o)=\left\{\begin{array}[]{ll}0&v_{\mathrm{pred}}\neq v_{\mathrm{gt}}\\ 1&v_{\mathrm{pred}}=v_{\mathrm{gt}}\\ 1&v_{\mathrm{pred}}=v_{\mathrm{variant}}\end{array}\right$. (5)
@@
where $v_{pred}$ represents the predicted verb in the output sentence $o$, and $v_{gt}$ represents the verb in the ground truth sentence. A full reward is obtained if they are equal, or if $v_{pred}$ is a variant $v_{variant}$ of $v_{gt}$, the same full reward is given too.
@@
Action Recognition. Action Recognition task trains and preserves the model’s action perception capability, and the output of the model is fixed to a single verb. We feed the model the ground truth video segment with prompt as “Use a verb to describe the event based on the video”, then compare the predicted verb against any verbs appearing in the ground truth query description. If the model’s predicted verb is present in the reference sentence, it receives a full reward. For example, if the model outputs “walk” and the ground-truth sentence is “A person walks away and laughs”, a full reward is granted because “walk” occurs in the reference. As before, verb in different tenses are treated as equivalent. The reward function is as follows:
@@
$\mathrm{r}_{\mathrm{AR}}(o)=\left\{\begin{array}[]{ll}0&v_{\mathrm{pred}}\notin S_{\mathrm{gt}}\\ 1&v_{\mathrm{pred}}\in S_{\mathrm{gt}}\\ 1&v_{\mathrm{variant}}\in S_{\mathrm{gt}}\end{array}\right$. (6)
where $v_{pred}$ represents the predicted verb $o$, and $S_{gt}$ is the set of verbs in the ground truth sentence. A full reward is given if $v_{pred}$ is present in the set of verbs in $S_{gt}$, or if a variant $v_{variant}$ of $v_{pred}$ exists within the set of verbs in $S_{gt}$.
Video Description. Video Description task is employed to train and maintain the model’s holistic perception of events, yielding a complete segment level description. Specifically, we feed the model the ground-truth video segment and prompt as “Describe what people have done based on the video”. A full reward is granted as long as ground-truth verbs appear in the output sentence of the model. For instance, if the ground-truth verb is “jump” and the model produces “A person jumps and laughs” the reward is awarded because “jump” is present. The reward function is as follows:
$\mathrm{r}_{\mathrm{VD}}(o)=\left\{\begin{array}[]{ll}0&v_{\mathrm{gt}}\notin S_{\mathrm{pred}}\\ 1&v_{\mathrm{gt}}\in S_{\mathrm{pred}}\\ 1&v_{\mathrm{variant}}\in S_{\mathrm{pred}}\end{array}\right$. (7)
@@
where $v_{gt}$ represents the ground truth verb, and $S_{pred}$ denotes the set of verbs in the sentence $o$ predicted by the model.
@@
Figure 3: Overview of the proposed Invert4TVG framework. During reinforcement learning training the Invert4TVG prevents conflicts between TVG and Invert-TVG tasks and ensures each can be optimized with independent gradients. The LVLM dynamically chooses between TVG tasks and Invert-TVG tasks according to different probabilities. Whenever an Invert-TVG task is selected, one of the three variants VC, AR or VD is chosen with equal probability.
IoU and Format Reward Functions

### v2

# Invert4TVG: A Temporal Video Grounding Framework with Inversion Tasks Preserving Action Understanding Ability

@@ Abstract
Temporal Video Grounding (TVG) aims to localize video segments corresponding to a given textual query, which often describes human actions. However, we observe that current methods, usually optimizing for high temporal Intersection-over-Union (IoU), frequently struggle to accurately recognize or understand the underlying actions in both the video and query, thus reducing the effectiveness of these methods. To address this, we propose a novel TVG framework that integrates inversion-based TVG as auxiliary objectives to maintain the model’s action understanding ability. We introduce three kinds of inversion TVG tasks derived from the original TVG annotations: (1) Verb Completion, predicting masked verbs (actions) in queries given video segments; (2) Action Recognition, identifying query-described actions; and (3) Video Description, generating descriptions containing query-relevant actions given video segments. These inversion tasks are entirely derived from the original TVG tasks and are probabilistically integrated with them within a reinforcement learning framework. By leveraging carefully designed reward functions, the model preserves its ability to understand actions, thereby improving the accuracy of temporal grounding. Experiments show our method outperforms state-of-the-art approaches, achieving a 7.1% improvement in R1@0.7 on Charades-STA for a 3B model.
@@
Figure 1: Left side: A specific example of temporal video grounding. According to the model’s reasoning process,it can be seen that our method achieves better understanding of actions in the video compared to VideoChat-R1 and Time-R1. Right side: statistical results demonstrating that Time-R1, which is optimized solely for the IoU loss, reduces action understanding accuracy (where VC, AR, and VD are the proposed three auxiliary inversion TVG tasks measuring multi-granularity action understanding ability). By introducing Inversion-TVG tasks, our method preserves action understanding ability and thus boosts TVG ability (as shown in R1@0.3, R1@0.5, and R1@0.7). Baseline is Qwen-2.5-VL-3B.
1 Introduction
@@
Temporal Video Grounding (TVG) is crucial for long-form video understanding (Gaidon et al., 2013; Laptev and Pérez, 2007; Darrell and Pentland, 1993). It localizes a video segment matching a textual query (Gao et al., 2017; Zhang et al., 2023), enabling applications like video-text retrieval (Zhang et al., 2024) and UAV positioning (Ju et al., 2024).
@@
Existing TVG approaches fall into three paradigms: (1) Traditional methods using hand-crafted features, sliding windows, and DETR-like networks (Shi et al., 2022; Gordeev et al., 2024); (2) LVLMs (Bai et al., 2025; Li et al., 2023) that regress segment duration via pretraining; and (3) RL-finetuned LVLMs, such as Time-R1 (Wang et al., 2025a), using Reinforcement Learning (RL) with format rewards for structured reasoning and IoU rewards for alignment.
Although significant progress has been made, we still find wrong grounding results in existing SOTA methods, and most of these wrong cases stem from incorrect action understanding. Figure 1 left shows one case. In the video, a man unbuttons his shirt, takes it off, puts on another one, and then buttons it up. The query is “A person putting on clothes and fastening the buttons”, which requires localizing the actions “putting” and “fastening”. Both VideoChat-R1 and Time-R1 notice a hand touching a button and localize the action as “buttoning” rather than “unbuttoning”, indicating that they seem to focus only on the button itself without distinguishing between buttoning and unbuttoning. We conjecture that these wrong groundings occur because video grounding models are generally optimized only for IoU. Although IoU is improved, this comes at the cost of reduced action understanding capability, which in turn limits their overall video grounding performance. Figure 1 right demonstrates this statistically. Time-R1 (Wang et al., 2025a), optimized for IoU, shows improvement on TVG metrics (e.g., R1@0.3/R1@0.5/R1@0.7) compared to the baseline Qwen2.5-VL-3B. However, it exhibits degradation in action understanding tasks (VC/VD/AR), which ultimately hampers its TVG accuracy.
A key insight of this work is that training a TVG model effectively requires jointly training auxiliary tasks to preserve the model’s action understanding capability. A naive approach to achieving this would be to train the TVG task alongside general action understanding tasks such as action recognition/detection/classification. However, these general tasks are not specifically designed for the temporal video grounding objective, and the understanding learned from them may not align well with the precise temporal localization required in TVG.
Unlike general action understanding tasks, we design action understanding tasks that are specifically tailored for the temporal video grounding task. Specifically, by inverting the input and output of the original TVG task, we convert the localization task into understanding task, obtaining a set of Invert-TVG tasks. A key advantage of these Invert-TVG tasks, compared to general action understanding tasks, is that they share the same training data as the original TVG task. On the same video-query data, our method performs both video localization (via the original TVG task) and action understanding (via the Invert-TVG tasks). This tight coupling enables the learned action understanding to be directly aligned with and supportive of the temporal grounding objective, resulting in more effective and synergistic learning.
@@
Specifically, given a video and a natural language query, the original TVG task predicts the temporal segment duration where the action occurs. Inversely, given a video segment, the proposed Invert-TVG tasks infer the action-related information defined in the query from the given segment. We introduce three Invert-TVG tasks: (1) Verb Completion (VC): mask verbs (actions) in the query and then infer the verbs from video segments. (2) Action Recognition (AR): classify the actions in a given video segment where the ground-truth action is in the query. (3) Video Description (VD): generate descriptions for a given video segment, and the descriptions should contain actions provided in the query.
With the well-defined Invert-TVG tasks, we then propose a reinforcement learning framework that optimizes TVG and Invert-TVG tasks together. However, for large-scale models (e.g., 3B/7B parameters), simultaneously optimizing multiple objectives incurs substantial memory overhead. Moreover, the TVG and Invert-TVG tasks are conflict: the ground-truth video segment that TVG is required to produce is precisely the input to an Invert-TVG task, and the original query that an Invert-TVG task may ask for is exactly the input to the TVG task. To address this, we adopt an alternating optimization strategy, executing TVG and Invert-TVG tasks interleavingly. Besides, since temporal video grounding is our main objective while action understanding is an auxiliary objective, we optimize the TVG task with a higher probability while using a lower probability for the Invert-TVG tasks.
Our contributions include:
@@
  • We identify action understanding degradation in TVG from IoU over-optimization, and address the problem via inversion TVG tasks.
@@
  • We design three inversion TVG tasks which are self-supervised tasks re-purposing TVG annotations for action understanding, including verb completion, action recognition, and video description.
@@
  • We propose a reinforcement learning framework dynamically balancing TVG and Invert-TVG tasks, ensuring robust grounding and understanding.
@@
Temporal Video Grounding. Temporal Video Grounding (TVG) (Gao et al., 2017; Hendricks et al., 2017) localizes specific segments in untrimmed videos based on natural language queries. Recent methods fall into two categories: feature-based and frame-based LVLM approaches. Feature-based methods (Carreira and Zisserman, 2017; Lin et al., 2022) extract video and text features using pretrained encoders, then predict timestamps via multimodal fusion. These rely heavily on feature quality, limiting performance. Frame-based LVLM methods have recently gained traction for their strong generalization capabilities.For instance, NumPro (Wu et al., 2025) introduces a frame-numbering mechanism akin to flipping a manga for efficient temporal grounding, while TimeSuite (Zeng et al., 2024) employs grounded tuning to enhance Large Language Models (LLMs) for long-form video understanding. While methods like these and others (Li et al., 2024; Ren et al., 2024) utilize supervised fine-tuning to generate event sequences, they can still struggle with precise boundary detection on benchmarks like Charades-STA compared to specialized feature-based approaches. To address this, Time-R1 (Wang et al., 2025a) employs reinforcement learning (RL) with IoU rewards, achieving state-of-the-art TVG performance. However, its focus on temporal metrics neglects semantic alignment, constraining long-form video understanding
@@
RL in LVLMs. RL has advanced post-training of LVLMs through Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022; Yu et al., 2024) and Reinforcement Learning with Verifiable Reward (RLVR) (DeepSeek-AI, 2025; Chen et al., 2025). RLHF aligns models with human preferences, improving tasks like image captioning, while RLVR enhances deterministic tasks like visual grounding (Liu et al., 2025). However, RL applications in long-form video tasks remain underexplored due to temporal complexity and semantic challenges. TimeZero (Wang et al., 2025b), Time-R1 (Wang et al., 2025a), VideoChat-R1 (Li et al., 2025) apply RL to TVG but overlooks semantic understanding degradation from IoU-focused rewards. Our Invert4TVG framework addresses this by repurposing TVG data into self-supervised tasks, enhancing action semantic understanding and surpassing traditional RL limitations in video grounding.
2 Method
@@
Figure 2: We propose three Invert-TVG tasks. By partially reversing the inputs and outputs of the TVG task we obtain Verb Completion, Action Recognition and Video Description, which reuse the original TVG dataset by taking ground truth video segments as input to reconstruct the target query related actions. The prompts for the three invert-TVG tasks are not identical. For VC, the verb in the query is removed, and the model is required to complete and fill in this verb. AR asks the model to directly estimate the verb in the video. VD requires the model to describe the video content containing action verbs in the query.
@@
The TVG task aims to temporally localize video segments within long-form videos based on natural language queries. Given a video $V$, and a language query $q$, the goal is to identify the temporal boundaries $\tau=[t_{s},t_{e}]$ of the segment of $V$ that best corresponds to $q$, where $t_{s},t_{e}\in\mathbb{R}^{+}$. The formal definition of the TVG task is as follows:
@@
Our approach is fundamentally a reinforcement learning algorithm that fine-tunes LVLMs (specifically the Qwen2.5-VL model series) by integrating both TVG and Invert-TVG tasks. In the following, we first introduce the fundamentals of GRPO (i.e., Group Relative Policy Optimization, a reinforcement learning algorithm proposed in (DeepSeek-AI, 2025)). Next, we introduce the proposed inversion TVG tasks, together with reward functions used to train the TVG and Invert-TVG tasks. Finally, we introduce our Invert4TVG reinforcement learning framework.
2.1 Preliminary of Group Relative Policy Optimization
DeepSeek-R1 (DeepSeek-AI, 2025), an early R1-style open-source LLM, uses GRPO to train policy $\pi_{\theta}$ for reasoning before answers. For query $q$, it generates responses $o_{1},\dots,o_{G}$} with score with $r(\cdot)$, and maximizes:
@@ 2.1 Preliminary of Group Relative Policy Optimization
2.2 Invert-TVG Tasks and Reward Functions
@@
In temporal video grounding, the accuracy of model inference highly depends on its understanding of both the video $V$ and the query $q$. Therefore, we do not rely solely on the IoU reward but also introduce additional rewards to keep or even enhance the model’s action understanding ability. To this end, as illustrated in Figure 2, we design three Invert-TVG tasks and define reward functions measuring how these tasks are fulfilled. Thanks to advances in NLP, a wealth of mature linguistic toolkits (e.g., SpaCy) can now effortlessly convert verbs into their various tenses or even bring them back to the root form, making it feasible for us to compute reward values for the following inversion tasks.
@@
Verb Completion (fine granularity). Verb Completion is a task that masks verbs in the query and asks the model to recover the verbs from the ground truth video segment. For example, if the original query is “Person closed the door”, the masked sentence is “Person [ ] the door”. The prompt is “Add a verb for the event ‘Person [ ] the door’ based on the video”. As long as the model outputs a sentence successfully recovering the verbs of the ground truth, a reward is given. For instance, if the output sentence is “Person [closes] the door” a full reward is given. Due to the randomness in the model’s output, we are primarily concerned with whether the model comprehends the actions within the relevant segments. Therefore, we treat verbs in different tenses as equivalent using SpaCy. The reward function is as follows:
$\mathrm{r}_{\mathrm{VC}}(o)=\left\{\begin{array}[]{ll}0&\mathrm{SpaCy}(v_{\mathrm{pred}})\neq\mathrm{SpaCy}(v_{\mathrm{gt}})\\ 1&\mathrm{SpaCy}(v_{\mathrm{pred}})=\mathrm{SpaCy}(v_{\mathrm{gt}})\end{array}\right$. (5)
@@
where $v_{pred}$ represents the predicted verb in the output sentence $o$, and $v_{gt}$ represents the verb in the ground truth sentence. A full reward is obtained if the root form of the verbs are equal, where $\mathrm{SpaCy}(\cdot)$ brings a verb to its root form.
@@
Action Recognition (middle granularity). Action Recognition task trains and preserves the model’s action perception capability, and the output of the model is fixed to a single verb. We feed the model the ground truth video segment with prompt as “Use a verb to describe the event based on the video”, then compare the predicted verb against any verbs appearing in the ground truth query description. If the model’s predicted verb is present in the reference sentence, it receives a full reward. For example, if the model outputs “walk” and the ground-truth sentence is “A person walks away and laughs”, a full reward is granted because “walk” occurs in the reference. As before, verb in different tenses are treated as equivalent. The reward function is as follows:
@@
$\mathrm{r}_{\mathrm{AR}}(o)=\left\{\begin{array}[]{ll}0&\mathrm{SpaCy}(v_{\mathrm{pred}})\notin S_{\mathrm{gt}}\\ 1&\mathrm{SpaCy}(v_{\mathrm{pred}})\in S_{\mathrm{gt}}\end{array}\right$. (6)
@@
where $v_{pred}$ represents the predicted verb $o$, and $S_{gt}$ is the set of root-formed verbs in the ground truth query.
Video Description (coarse granularity). Video Description task is employed to train and maintain the model’s holistic perception of events, yielding a complete segment level description. Specifically, we feed the model the ground-truth video segment and prompt as “Describe what people have done based on the video”. A full reward is granted as long as ground-truth verbs appear in the output sentence of the model. For instance, if the ground-truth verb is “jump” and the model produces “A person jumps and laughs” the reward is awarded because “jump” is present. The reward function is as follows:
$\mathrm{r}_{\mathrm{VD}}(o)=\left\{\begin{array}[]{ll}0&\mathrm{SpaCy}(v_{\mathrm{gt}})\notin S_{\mathrm{pred}}\\ 1&\mathrm{SpaCy}(v_{\mathrm{gt}})\in S_{\mathrm{pred}}\end{array}\right$. (7)
where $v_{gt}$ represents the ground truth verb, and $S_{\mathrm{pred}}$ denotes the set of root-formed verbs in the sentence $o$ predicted by the model.
@@
Figure 3: Overview of the proposed Invert4TVG framework. The LVLM dynamically chooses between TVG tasks and Invert-TVG tasks according to different probabilities. Whenever an Invert-TVG task is selected, one of the three variants VC, AR or VD is chosen with equal probability.
2.3 IoU and Format Reward Functions